# General effective reduction theory of integral polynomials of given non-zero discriminant and its applications

### v1

@@ B. Generalizations to rationally monogenic orders
From work of Delone and Faddeev (1940) it follows that $RM(3)\leq 1$, that is, every order of a cubic number field is at most one time rationally monogenic. From a result of Bérczes, Evertse and Győry (2004) the following analogue of Theorem A.1 can be deduced:
Theorem B.2.
@@
Similarly to Theorem A.2 the proof uses the result of Beukers and Schlickewei (1996) mentioned above.
@@
The proof of part (ii) is similar to that of Theorem B.2 but with a combinatorial improvement in the argument. The proof of part (i) also uses a parametrization of quartic rings and their cubic resolvent rings.
Recently, the following analogue of Theorem A.4 for rationally monogenic orders was proved:
Theorem B.4 (Evertse, 2023).
@@
Part (i) is best possible in the sense that there are quartic number fields having infinitely many two times rationally monogenic orders. It is not clear whether the $5$-transitivity condition on the Galois closure of $K$ in part (ii) is necessary; this was just a technical condition needed for the proof.
Similary to Theorem A.4, Theorem B.4 has been proved by means of a reduction to unit equations in more than two unknowns, and a use of ineffective finiteness theorems for such equations. So likewise, Theorem B.4 is ineffective.
We would like to finish with a connection with Hermite equivalence classes, discussed in Subsection 3.2. Let $\alpha$ be an algebraic number of degree $n$. Let $\mathcal{I}_{\alpha}:=\mathbb{Z}_{\alpha}+\alpha\mathbb{Z}_{\alpha}$ be the fractional ideal of $\mathbb{Z}_{\alpha}$ generated by $1$ and $\alpha$. This is known to be invertible, see Simon (2003). It is called also the invariant ideal of $f_{\alpha}$.
Recall that a polynomial in $\mathbb{Z}[X]$ is primitive if its coefficients have greatest common divisor $1$.
Theorem B.5 (BEGyRS, 2023).
Let $f,g\in\mathbb{Z}[X]$ be two primitive, irreducible polynomials. Then the following two assertions are equivalent:
  • (i)
    $f$ and $g$ are Hermite equivalent;
  • (ii)
In the particular case that $f$ and $g$ are monic, we have $\alpha\in\mathbb{Z}[\alpha]=\mathbb{Z}_{\alpha}$ and $\mathcal{I}_{\alpha}=\mathbb{Z}_{\alpha}$ and likewise for $g$ and $\beta$. So two monic, irreducible polynomials $f,g\in\mathbb{Z}[X]$ are Hermite equivalent if and only if $f$ has a root $\alpha$ and $g$ a root $\beta$ such that $\mathbb{Z}[\alpha]=\mathbb{Z}[\beta]$.
Combining Theorems B.5 and B.4 one can deduce the following counterpart of Theorem 3.4. For a number field $K$, let $\mathcal{P}\mathcal{I}(K)$ denote the set of primitive, irreducible polynomials in $\mathbb{Z}[X]$ having a root $\alpha$ such that $K=\mathbb{Q}(\alpha)$.
Theorem B.6 (Evertse, 2023).
@@
In a later, yet to be written extended version of this paper, we will give more details and background about the above mentioned theorems.
C. Monogenicity, class group and Galois group
@@
Recently, it has been proved in a precise and quantitative form that the monogenicity has an increasing effect on the class groups of number fields and orders; see Bhargava and Varma (2016), Ho, Shankar and Varma (2018), Bhargava, Hanke and Shankar (2020), Siad (2021), Swaminathan (2023).
The examples of degree $n=3,4,5,6$ in Section 6 show that the multiplicity of monogenicity can be relatively large if the Galois group is $S_{n}$, i.e. if its size is large relative to $n$.
@@
Recently, Arpin, Bozlee, Herr and Smith (2023) studied so-called twisted monogenic relative extensions $K/L$. They proved that $L$ has trivial class group (this is the case if e.g. $L=\mathbb{Q}$) if and only if every twisted monogenic extension of $L$ is monogenic.
D. Distribution of monogenic number fields

### v2

@@ 10.3. Generalizations for rationally monogenic orders
From work of Delone and Faddeev (1940) it follows that $RM(3)\leq 1$, that is, every order of a cubic number field is at most one time rationally monogenic (and in fact precisely one time if the order is primitive). From a result of Bérczes, Evertse and Győry (2004) the following analogue of Theorem 10.3 can be deduced:
Theorem 10.13.
@@
Similarly to Theorem 10.3 the proof uses Theorem 10.6 of Beukers and Schlickewei (1996) mentioned above.
@@
The proof of part (ii) is similar to that of Theorem 10.13 but with a combinatorial improvement in the argument. The proof of part (i) also uses a parametrization of quartic rings and their cubic resolvent rings.
Recently, the following analogue of Theorem 10.5 for rationally monogenic orders was proved:
Theorem 10.15 (Evertse, 2023).
@@
Part (i) is best possible in the sense that there are quartic number fields having infinitely many two times rationally monogenic orders. In fact, Bérczes, Evertse and Győry (2013, end of Section 1) give the following construction:
Let $r,s$ be integers such that $f(X)=(X^{2}-r)^{2}-X-s$ is irreducible, and let $K=\mathbb{Q}(\alpha)$, where $\alpha$ is a root of $f$. Then $K$ has infinitely many orders $\mathcal{O}_{m}$ ($m=1,2,\ldots$) with the following property: $\mathcal{O}_{m}=\mathbb{Z}[\alpha_{m}]=\mathbb{Z}[\beta_{m}]$, where $\beta_{m}=\alpha_{m}^{2}-r_{m}$, $\alpha_{m}=\beta_{m}^{2}-s_{m}$ for some integers $r_{m},s_{m}$.
It is clear that $\alpha_{m},\beta_{m}$ in the above theorem are not ${\rm GL}_{2}(\mathbb{Z})$-equivalent. We would like to pose the following problem:
Problem 5.
Does every quartic number field have infinitely many orders that are two times rationally monogenic? If not, can we characterize those quartic number fields that do? Do the two times rationally monogenic orders have a particular structure?
Similary to Theorem 10.5, Theorem 10.15 has been proved by means of a reduction to unit equations in more than two unknowns, and a use of ineffective finiteness theorems for such equations. So likewise, Theorem 10.15 is ineffective.
It is not clear whether the $5$-transitivity condition on the Galois closure of $K$ in part (ii) is necessary; this was just a technical condition needed for the proof. We are interested in the following problem:
Problem 6.
Is it true that every number field of degree $n\geq 5$ has only finitely many orders that are two times rationally monogenic? If not, can we characterize those number fields that do?
Combining Theorems 5.11 and 10.15 one can deduce the following counterpart of Theorem 3.4. For a number field $K$, let $\mathcal{P}\mathcal{I}(K)$ denote the set of primitive, irreducible polynomials in $\mathbb{Z}[X]$ having a root $\alpha$ such that $K=\mathbb{Q}(\alpha)$.
Corollary 10.16 (Evertse, 2023).
@@
10.4. Outlines of the proofs of Theorems 10.14 (ii) and 10.15
a
The main new tool is the following result.
Theorem 10.17.
Let $F$ be a field of characteristic 0, and $\Gamma$ a finitely generated subgroup of $F^{*}$. Then there is a finite subset $S\SS$ of $F^{*}$ with $S1\in\SS$, such that for the set of solutions $(x_{1},x_{2},x_{3},y_{1},y_{2},y_{3})\in\Gamma^{6}$ of
(10.5) $(x_{1}-1)(x_{2}-1)(x_{3}-1)=(y_{1}-1)(y_{2}-1)(y_{3}-1)$
at least one of the following holds:
(i) at least one of $x_{1},\ldots,y_{3}$ belongs to $\mathcal{S}$;
(ii) there are $\eta_{1},\eta_{2},\eta_{3}\in\{\pm 1\}$ such that $(y_{1},y_{2},y_{3})$ is a permutation of $(x_{1}^{\eta_{1}},x_{2}^{\eta_{2}},x_{3}^{\eta_{3}})$;
(iii) one of the numbers in $\{x_{i}x_{j},\,x_{i}/x_{j},\,y_{i}y_{j},\,y_{i}/y_{j}:\,1\leq i<j\leq 3\}$ is equal to either 1-1, or to a primitive cube root of unity.
Proof.
This is Proposition 8.1 of Bérczes, Evertse and Győry (2013). The proof is basically to expand (10.5), divide by one term to get an equation of type (10.4) in $16$ terms equal to $1$, consider all possible partitions into minimal vanishing subsums, and apply Theorem 10.8 to each of them (by using symmetric properties we can substantially reduce the number of cases). ∎
We need some other lemmas. Let $K$ be a number field of degree $n\geq 4$. Denote by $G$ the normal closure of $K$ and by $x\mapsto x^{(i)}$ ($i=1,\ldots,n$) the embeddings of $K$ in $G$. For $\alpha$ with $\mathbb{Q}(\alpha)=K$ we define the cross ratios
${\rm cr}_{ijkl}(\alpha):=\frac{(\alpha^{(i)}-\alpha^{(j)})(\alpha^{(k)}-\alpha^{(l)})}{(\alpha^{(i)}-\alpha^{(k)})(\alpha^{(j)}-\alpha^{(l)})}$
for any four distinct indices $i,j,k,l\in\{1,\ldots,n\}$ and we define the tuple
$\lambda(\alpha):=({\rm cr}_{123i}(\alpha),{\rm cr}_{1i32}(\alpha):\,i=4,\ldots,n).$
We call $\alpha,\beta\in K$ $GL_{2}(\mathbb{Q})$-equivalent if $\beta=\mbox{\large{$\textstyle{\frac{a\alpha+b}{c\alpha+d}}$}}$ for some $\big(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big)\in GL_{2}(\mathbb{Q})$.
Lemma 10.18.
Let $\alpha,\beta$ with $\mathbb{Q}(\alpha)=\mathbb{Q}(\beta)=K$.
(i) $\lambda(\alpha)=\lambda(\beta)$ $\Longleftrightarrow$ $\alpha,\beta$ are $GL_{2}(\mathbb{Q})$-equivalent.
(ii) If $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$ and $\alpha,\beta$ are $GL_{2}(\mathbb{Q})$-equivalent, then $\alpha,\beta$ are $GL_{2}(\mathbb{Z})$-equivalent.
Proof.
[…]
(ii). See for instance Lemma 2.6 of Evertse (2023). ∎
Lemma 10.19.
Let $\alpha,\beta$ with $\mathbb{Q}(\alpha)=\mathbb{Q}(\beta)=K$ and $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$. Then for all distinct $i,j,k,l\in\{1,\ldots,n\}$ we have ${\rm cr}_{ijkl}(\beta)/{\rm cr}_{ijkl}(\alpha)\in\mathcal{O}_{G}^{*}$.
Proof.
This is Lemma 2.4 of Evertse (2023). ∎
Sketch of the proof of Theorem 10.14 (ii).
Let $\mathcal{O}$ be an order of $K$. Note that for $\alpha\in K$ with $K=\mathbb{Q}(\alpha)$ we have relations
${\rm cr}_{123i}(\alpha)+{\rm cr}_{1i32}(\alpha)=1\ \ (i=4,\ldots,n).$
By Lemma 17.7.3 of Evertse and Győry (2017), the set of tuples
$\{\lambda(\alpha):\,\mathbb{Z}_{\alpha}=\mathcal{O}\}$
generates a multiplicative subgroup of $(G^{*})^{2n-6}$ of rank at most n(n1)/2n(n-1)/2. In the deduction of this we used a refinement of Lemma 10.19. Now an application of Corollary 10.7 and Lemma 10.18 implies Theorem 10.14 (ii). ∎
Call $\alpha_{1}$ with $\mathbb{Q}(\alpha_{1})=K$ $k$-special if there are $\alpha_{2},\ldots,\alpha_{k}\in K$ such that $\alpha_{1},\ldots,\alpha_{k}$ are pairwise $GL_{2}(\mathbb{Z})$-inequivalent and $\mathbb{Z}_{\alpha_{1}}=\cdots=\mathbb{Z}_{\alpha_{k}}$. We should mention here that if $K$ has degree $3$ then there are no $2$-special numbers in $K$.
In the proof of Theorem 10.15, we need the following lemma.
Lemma 10.20.
Assume $n\geq 4$. Let $\mathcal{C}$ be a $GL_{2}(\mathbb{Q})$-equivalence class of $2$-special numbers. Then $\mathcal{C}$ is the union of finitely many $GL_{2}(\mathbb{Z})$-equivalence classes.
Idea of proof.
This is Proposition 5.1 of Evertse (2023). Its proof is fairly complicated. We give a brief outline.
We define ${\rm cr}_{ijkl}(\mathcal{C}):={\rm cr}_{ijkl}(\alpha)$ for any $\alpha\in\mathcal{C}$. This is well-defined since $GL_{2}(\mathbb{Q})$-equivalent algebraic numbers have the same cross ratios. Let $\alpha\in\mathcal{C}$, let $\beta\in K$ be such that $\mathbb{Z}_{\beta}=\mathbb{Z}_{\alpha}$ and $\beta$ is not $GL_{2}(\mathbb{Z})$-equivalent to $\alpha$, and let $\mathcal{D}$ be the $GL_{2}(\mathbb{Q})$-equivalence class of $\beta$. Then $\mathcal{D}\not=\mathcal{C}$ by Lemma 10.18 (ii). Clearly, ${\rm cr}_{ijkl}(\beta)=:{\rm cr}_{ijkl}(\mathcal{D})$ depends only on $\mathcal{D}$. By Lemma 10.19 we have ${\rm cr}_{ijkl}(\mathcal{D})/{\rm cr}_{ijkl}(\mathcal{C})\in\mathcal{O}_{G}^{*}$ for all $i,j,k,l$. Further,
$1={\rm cr}_{ijkl}(\beta)+{\rm cr}_{ilkj}(\beta)={\rm cr}_{ijkl}(\mathcal{C})\cdot\frac{{\rm cr}_{ijkl}(\mathcal{D})}{{\rm cr}_{ijkl}(\mathcal{C})}+{\rm cr}_{ilkj}(\mathcal{C})\cdot\frac{{\rm cr}_{ilkj}(\mathcal{D})}{{\rm cr}_{ilkj}(\mathcal{C})}$
for all $i,j,k,l$. Now by Theorem 10.6, for given $\mathcal{C}$ there are only finitely many possible values for each ${\rm cr}_{ijkl}(\mathcal{D})$ and thus, by Lemma 10.18, at most finitely many possibilities for $\mathcal{D}$. It follows that $\mathcal{C}$ is the union of finitely many sets
$\mathcal{C}(\mathcal{D}):=\{\alpha\in\mathcal{C}:\,\text{there is $\beta\in\mathcal{D}$ with $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$}\}$
where $\mathcal{D}\not=\mathcal{C}$ is a $GL_{2}(\mathbb{Q})$-equivalence class of $2$-special numbers. So it suffices to prove that each set $\mathcal{C}(\mathcal{D})$ is the union of finitely many $GL_{2}(\mathbb{Z})$-equivalence classes.
Now fix $\mathcal{D}$, $\alpha\in\mathcal{C}(\mathcal{D})$, and $\beta\in\mathcal{D}$ such that $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$. Let $\alpha^{\prime}$ be any other element of $\mathcal{C}(\mathcal{D})$. Then we can write
$\alpha^{\prime}=\frac{a\alpha+b}{c\alpha+d}\ \text{with }a,b,c,d\in\mathbb{Z},\,{\rm gcd}(a,b,c,d)=1,\ ad-bc=:\Delta\not=0.$
We have to prove that the numbers $\alpha^{\prime}\in\mathcal{C}(\mathcal{D})$ lie in only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes. Recall that there are $U\in GL_{2}(\mathbb{Z})$ and $a^{\prime},b^{\prime},d^{\prime}\in\mathbb{Z}$ with $a^{\prime}d^{\prime}=\Delta$, $|b^{\prime}|\leq|d^{\prime}|/2$ such that $U\big(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big)=\big(\begin{smallmatrix}a^{\prime}&b^{\prime}\\ 0&d^{\prime}\end{smallmatrix}\big)$. Hence $\alpha^{\prime}$ is $GL_{2}(\mathbb{Z})$-equivalent to $\alpha^{*}:=(a^{\prime}\alpha+b^{\prime})/d^{\prime}$. It suffices to prove that there are only finitely many possibilities for $\alpha^{*}$. It is in fact sufficient to prove that $\Delta$ is bounded, since for given $\Delta$ there are only finitely many possibilities for $(a^{\prime},b^{\prime},d^{\prime})$. The boundedness of $\Delta$ is provided by the following elementary lemma, which is Proposition 4.1 of Evertse (2023). We refer to that paper for the rather lengthy proof.
Lemma 10.21.
Let $D$ be the discriminant of $\mathbb{Z}_{\alpha}$, and let $\mathfrak{a}(\alpha,\beta)$ be the ideal of $\mathcal{O}_{G}$ generated by the numbers ${\rm cr}_{ijkl}(\beta)/{\rm cr}_{ijkl}(\alpha)-1$ ($1\leq i<j<k<l\leq n$). Then $\Delta$ divides $D^{5}\cdot\mathfrak{a}(\alpha,\beta)^{2}$.
Sketch of the proof of Theorem 10.15.
For $\alpha,\beta$ with $\mathbb{Q}(\alpha)=\mathbb{Q}(\beta)=K$ and distinct $i,j,k,l\in\{1,\ldots,n\}$ we put
$\varepsilon_{ijkl}(\alpha,\beta):=\frac{{\rm cr}_{ijkl}(\beta)}{{\rm cr}_{ijkl}(\alpha)}.$
Lemma 10.19 implies that if $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$, then $\varepsilon_{ijkl}(\alpha,\beta)\in\mathcal{O}_{G}^{*}$.
The case $n=4$. We have to show that there are only finitely many orders $\mathcal{O}$ of $K$ such that $\mathcal{O}=\mathbb{Z}_{\alpha}$ for some $3$-special $\alpha$. It clearly suffices to show that the $3$-special numbers $\alpha$ lie in only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes.
Let $\alpha\in K$ be $3$-special, and choose $\beta,\gamma\in K$ such that $\alpha,\beta,\gamma$ are pairwise $GL_{2}(\mathbb{Z})$-inequivalent, and $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}=\mathbb{Z}_{\gamma}$. Let $(i,j,k,l)$ be a permutation of $(1,2,3,4)$. Then the equation
${\rm cr}_{ijkl}(\alpha)x+{\rm cr}_{ilkj}(\alpha)y=1$
has three distinct solutions $(x,y)\in\mathcal{O}_{G}^{*}\times\mathcal{O}_{G}^{*}$, i.e., $(1,1)$, $(\varepsilon_{ijkl}(\alpha,\beta),\varepsilon_{ilkj}(\alpha,\beta))$, $(\varepsilon_{ijkl}(\alpha,\gamma),\varepsilon_{ilkj}(\alpha,\gamma))$. Now Theorem 10.9 implies that ${\rm cr}_{ijkl}(\alpha)$ can assume only finitely many values. From Lemma 10.19 it now follows that the $3$-special $\alpha\in K$ lie in only finitely $GL_{2}(\mathbb{Q})$-equivalence classes. Finally, from Lemma 10.20 it follows that they lie in finitely many $GL_{2}(\mathbb{Z})$-equivalence classes.
The case $n\geq 5$. We have to show that there are only finitely many orders $\mathcal{O}$ of $K$ such that $\mathcal{O}=\mathbb{Z}_{\alpha}$ for some $2$-special $\alpha$. It clearly suffices to show that the $2$-special numbers $\alpha$ lie in only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes.
Let $\alpha\in K$ be $2$-special, and choose $\beta$ such that $\alpha,\beta$ are $GL_{2}(\mathbb{Z})$-inequivalent and $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$. Henceforth, we write $\varepsilon_{ijkl}$ for $\varepsilon_{ijkl}(\alpha,\beta)$. Let $i,j,k,l$ be distinct indices from $\{1,\ldots,n\}$. Then
${\rm cr}_{ijkl}(\alpha)+{\rm cr}_{ilkj}(\alpha)=1,\ \ {\rm cr}_{ijkl}(\alpha)\varepsilon_{ijkl}+{\rm cr}_{ilkj}(\alpha)\varepsilon_{ilkj}=1,$
$\varepsilon_{ilkj}/\varepsilon_{ijkl}=\varepsilon_{iljk}$, which imply
(10.6) ${\rm cr}_{ijkl}(\alpha)=\frac{\varepsilon_{ilkj}-1}{\varepsilon_{ilkj}-\varepsilon_{ijkl}},\ \ {\rm cr}_{ijkl}(\beta)=\varepsilon_{ijkl}{\rm cr}_{ijkl}(\alpha)=\frac{\varepsilon_{ilkj}-1}{\varepsilon_{iljk}-1}.$
Now picking a fifth index $m$, and using $\mbox{\large{$\textstyle{\frac{{\rm cr}_{jmlk}(\beta){\rm cr}_{ijkm}(\beta)}{{\rm cr}_{ijkl}(\beta)}}$}}=1$, we obtain
(10.7) $\frac{\varepsilon_{jklm}-1}{\varepsilon_{jkml}-1}\cdot\frac{\varepsilon_{imkj}-1}{\varepsilon_{imjk}-1}\cdot\frac{\varepsilon_{iljk}-1}{\varepsilon_{ilkj}-1}=1.$
We apply Theorem 10.17 to (10.7) for all $i,j,k,l,m$. Our assumption that the Galois group of $G$ is $5$-transitive implies various conjugacy relations between the $\varepsilon_{ijkl}$. Using all of these, we infer that for each quadruple $i,j,k,l$ there are only finitely many possible values for $\varepsilon_{ijkl}$ (we should mention here that without the $5$-transitivity assumption, we do not know how to prove this). Now (10.6) implies that there are only finitely many possible values for ${\rm cr}_{ijkl}(\alpha)$, if $\alpha$ runs through the $2$-special numbers of $\alpha$, and thus, by Lemma 10.18, that the $2$-special $\alpha\in K$ lie in only finitely many $GL_{2}(\mathbb{Q})$-equivalence classes. Finally, from Lemma 10.20 it follows that they lie in only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes. ∎
Appendix: Related topics
We briefly discuss some further topics related to monogenic number fields and monogenic orders and generalizations thereof that do not strictly belong to the reduction theory of integral polynomials.
A. Monogenicity, class group and Galois group
@@
Recently, surprising results have been obtained in precise and quantitative form that imply that on average, the monogenicity of a number field has an altering effect on the structure of its $2$-class group, see Bhargava, Hanke and Shankar (2020), Siad (2021), Swaminathan (2023), Shankar, Siad and Swaminathan (2025), and Bhargava, Shankar and Swaminathan (2025). The $2$-class group $Cl_{2}(K)$ of a number field $K$ is the group of ideal classes of $K$ whose order divides $2$.
To illustrate this, we recall some results from the literature. A monogenized number field is a pair $(K,\alpha)$ consisting of a number field $K$ and $\alpha\in\mathcal{O}_{K}$ such that $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$. Two monogenized number fields $(K_{1},\alpha_{1})$, $(K_{2},\alpha_{2})$ are called isomorphic if there are a field isomorphism $\varphi:\,K_{1}\to K_{2}$ and a rational integer $a$ such that $\alpha_{2}=\pm\varphi(\alpha_{1})+a$.
We now restrict ourselves to monogenized cubic fields. We define the height of a monogenized cubic field $(K,\alpha)$ as follows. Let $f=X^{3}+aX^{2}+bX+c\in\mathbb{Z}[X]$ be the minimal polynomial of $\alpha$. Then the height of $(K,\alpha)$ is
$H(K,\alpha):=$ $\max(|I(f)|^{3},J(f)^{2}/4),$
$\quad\text{where }I(f):=a^{2}-3b,\ \ J(f):=-2a^{3}+9ab-27c.$
One can show that isomorphic monogenized cubic fields have the same height. Further, the pair $(I(f),J(f)^{2})$ uniquely determines an isomorphism class. Lastly, the discriminant of $f$ is $D(f)=\mbox{\large{$\textstyle{\frac{1}{27}}$}}\big(4I(f)^{3}-J(f)^{2}\big)$.
Theorem A.1 (Bhargava, 2005).
Let $K$ run through the cubic number fields, ordered by discriminant.
(i) The average size of $Cl_{2}(K)$ over the totally real cubic fields is $5/4$.
(ii) The average size of $Cl_{2}(K)$ over the complex cubic fields is $3/2$.
Theorem A.2 (Bhargava, Hanke and Shankar, 2020).
Let $(K,\alpha)$ run through the monogenized cubic number fields whose Galois closure has Galois group isomorphic to $S_{3}$, ordered by height.
(i) The average size of $Cl_{2}(K)$ over the totally real monogenized cubic fields is $3/2$.
(ii) The average size of $Cl_{2}(K)$ over the complex monogenized cubic fields is $2$.
Siad (2021) proved a generalization of the last theorem for number fields of odd degree $\geq 5$.
@@
We briefly discuss some other topics. Recently, Arpin, Bozlee, Herr and Smith (2023a,b) introduced and studied twisted monogenic relative extensions $K/L$. They proved that $L$ has trivial class group (this is the case if e.g. $L=\mathbb{Q}$) if and only if every twisted monogenic extension of $L$ is monogenic.
Another topic worth of study is the connection between (multiplicity of) monogenicity of the ring of integers of a number field $K$ and the size of the Galois group of its Galois closure. The examples of number fields $K$ of degree $n=3,4,5,6$ in Section 6 show that the multiplicity of monogenicity of $\mathcal{O}_{K}$ can be relatively large if the Galois group of the Galois closure of $K$ is $S_{n}$, i.e. if its size is large relative to $n$.
B. Distribution of monogenic and non-monogenic number fields